OpenStudy (anonymous):

can some help me find the slope and y-intercept on 8x+7y=24

OpenStudy (anonymous):

Hint: Write down above equation in the form y=mx+c...where m is the slope and c as y-intercept.. So can you write 8x+7y=24 in the form y=mx+c?

OpenStudy (anonymous):

slope= -8/7 y int= 24/7

OpenStudy (petiteme):

Yup that's right! :)